### v1

# High stakes exams inflate gender gap thus leading to systematic grading errors in introductory physics

Abstract
Previous research has suggested that changing the percentage of the course grade associated with exam grades in STEM courses can change the gender gap in the course. It’s also been shown that assessments with the highest stakes have the lowest (relative) scores for female students. Previous research by the authors has shown that the implementation of retake exams can eliminate the gender gap in introductory physics courses. This paper explores several different hypotheses for why retake exams are associated with a zeroed gender gap. Analyzing data from exams with different stakes, we argue that the entire gender gap on introductory physics exams may be due to the stakes associated with those exams. In other words, the data support the idea that a gender grade-gap on exams is not measuring a gender difference in the physics knowledge or physics ability of these students.
preprint: APS/123-QED
I Introduction
In a recent paper [1] we showed data in support of the argument that demographic grade gaps are a result of course design (the Course Deficit model [2]) rather than the more common argument that the gaps result from average deficits in specific demographic groups of students (the Student Deficit model [3]). One way we did this was by showing that the gender gap disappeared in introductory physics classes that were changed to offer exam retake opportunities to all students, while the gender gap had been persistent in regular unchanged courses. Our previous work did not examine the mechanism for how the retakes eliminated the gender gap. For example, were women more likely to take advantage of the retake opportunity, or did women improve more than men on the second exam? In this paper we analyze more detailed assessment data on the gender gap issue in these courses and argue that a reasonable interpretation of our results is that the entire gender gap in these courses is determined by the stakes associated with the exams and does not reflect a gender difference in physics knowledge, opportunity, or ability.
I.1 Gender gap in physics
Gender equity gaps in introductory physics courses and exams have been studied extensively over the past several decades. Many different explanations for their existence have been proposed and investigated. Some of the leading theories for the gender gaps in physics are that women on average i) have poorer physics and/or math preparation [4, 5], ii) have less overall interest in physics [6], and/or iii) suffer more from test anxiety, which is amplified on high-stakes exams [7, 8, 9, 5, 10, 11, 12]. It’s also been shown that women don’t feel like they belong in physics as much as their male peers [13], that they have less of a science or physics identity [14], and/or have less self-efficacy in physics [5].
In a comprehensive review of gender gaps in concept inventories, Madsen et al. [15] found that none of the 30 different factors they considered fully explained the gender gap and thus concluded that there may be many interconnected issues leading to these gaps. Despite these different competing factors, there have been a few studies that have seemed to entirely eliminate the gender gap (example: [1]), suggesting that, despite the gap’s persistence and the uncertainty concerning its cause, it is possible to mitigate the effects giving rise to the gap.
Exploring one potentially fruitful mitigation, Cotner and Ballen [2] show that exam gender gaps in some biological science courses could be changed by lowering the stakes associated with exams. They compared overall success in different biology courses by categorizing them as either low-stakes (less than 40% of the course grade determined by quizzes and exams) or high stakes (more than 40% of the course grade determined by quizzes and exams). The men outperformed the women in the high-stakes courses, while there was no statistical difference between the genders in low-stakes courses.
Cotner et al. later suggest [7] that the reason for the change in performance when the stakes change is because women are more likely to suffer from anxiety which is amplified in high-stakes situations. In fact the connection between high-stakes exams and anxiety in women is well studied [7, 8, 9, 5, 10, 11, 12]. Maespina & Singh [5] found that women under-performed on high-stakes exams in calculus-based introductory physics courses. However, when they controlled for test anxiety, the gap became non-significant. There is evidence that test anxiety is an issue for women in other STEM majors as well. Espinosa et al. [10] found that in a large introductory biology course, women expressed higher test anxiety than men and that anxiety was a significant predictor for exam grades even when controlling for overall GPA. This is important not only in biology courses but also because a large population of introductory physics students are biology students. Similarly, Malesina et al. [9] find the same is true for bioscience students enrolled in a physics course. Salehi et al [8] study performance across three different assessment types, exams, non-exams and labs, in five different lower division STEM courses. They found that while there was a significant negative gap for women on exams, women consistently outperformed men on the other two types of assessments. Interestingly, they found that exam performance for women was negatively influenced by test anxiety in both lower and upper division courses, even though by the upper division courses the gender gap on exam scores had disappeared in both biology courses and other STEM courses.
While the evidence for anxiety makes it a compelling explanation for the gender gap, we instead propose that the high-stakes exams are the cause of the gender gap, and therefore it is the course structure that deserves mitigation. As an example for why we take this approach, consider a study performed by Simmons & Heckler [16] where they simulated new course grades for students (after a course had already been completed) by reducing the weight of the exams in the overall course grade. They found that making the exams worth less of the overall course grade, equity gaps in the course were reduced for women and other historically marginalized students compared to their non-marginalized peers. In this example, women’s anxiety was not changed, but instead the course structure was changed to successfully address the inequity. While this may seem like a semantics argument, positioning the stakes as the possible cause of the gender gap (instead of anxiety) focuses the research and intervention efforts on the changeable aspects of the course, rather than the (potentially unchangeable) aspects of entire populations. Similarly, it removes the urge to “fix” any perceived inadequacies among an underrepresented group (in this case, women).
I.2 Retake Exams in Physics
If high-stakes are the cause of the gender gap, one obvious solution would be to eliminate the exams as an assessment practice. Critics of this approach say that low-stakes assessments (like homework or group projects) are not as accurate as other forms of assessment because low-stakes assessments like homework allow students to use outside resources, collaboratively solve problems and potentially even find solutions they can copy without understanding. Others claim that there is value in measuring what a student can do without outside resources, and/or in a high pressure environment, which homework assignments won’t measure. While many, including Ref. [17] and also the authors of this paper, disagree with these critiques, if exams (timed solo tests or quizzes with little to no allowed resources) are measuring something important to many instructors, eliminating them entirely is not necessarily a viable option. If so, then mitigation of the harms/biases caused by these exams might be the best one can do.
One possible way to lower the stakes of an exam, while still using exams as an assessment is to offer a retake option for each exam. Thus far, there is very little research on retake exams in large introductory STEM classes. Davis et al. [18] find that students in a large introductory human physiology course who took advantage of retake exams on average increase their grade by 6%. Kortemeyer et al. [19] share the benefits of retakes as learning tools in introductory physics. However, none of the examples we found examined the relationship between the retakes and the gender gap in physics courses aside from our own previous work on this same dataset.
I.3 Models of achievement and assessment equity
This work, like our prior research [1] utilizes two theoretical models. We define a course as being equitable if student grades are not predictable by demographic, meaning that all demographic groups achieve the same overall grades on average. This is the definition of an Equity of Parity model of equity [20]. We also use Cotner & Ballen’s [2] Course Deficit Model. A Course Deficit Model assumes that any gaps in achievement within the course between any demographic groups are a product of the course structure. Importantly, this does not mean that inequities prior to the course do not exist, the course deficit model merely assumes that the equity gaps measured within the course are a product of the course itself. Our previous work [1] shows that by changing course structures it is possible to eliminate equity gaps and thus achieve equity of parity.
We have previously studied grading and assessment in this same physics course and have found that percent scale grading provides an additional penalty to underrepresented minority groups when compared to 4.0 scale grading [21, 22]. We also find [23] that on average, women leave slightly more blanks on exams than men, while men are more likely to miss entire quizzes. Given that there seem to be both structural biases and behavioral test-taking differences between different demographic groups, it seems prudent to examine the impact of certain assessment strategies on different demographics.
II Research questions
In this paper we make the assumption that because the gender gap decreased or was entirely eliminated in courses that use a retake exam option, that the reason for increased gender gaps in the intro physics courses in our data set is likely due to the assessment strategy in courses. This leads us to test a few different hypotheses:
  1. 1.
    Does the gender gap decrease in retake courses because students identifying as women are more likely to take advantage of the retake opportunity?
  2. 2.
    Does the gender gap decrease in retake courses because students identifying as women are doing better on their second try? (Presumably because the second try has allowed them to ‘catch up’.)
  3. 3.
    How are the stakes of an exam correlated with the gender gap?
III Database Details
All of the exam data we discuss here are from classes in the introductory physics series taken mostly by bioscience students and offered between 1997 and 2015 at an R1 public university. We have access to individual exam scores for three of the four classes that offered students retake exams. To compare the exam scores from the retake classes to the exam scores from the regular classes which did not offer retakes we add, to the database, exam data of the 254 non-retake classes from the same series for which we have scores from exams given to students during the term. We also added each student’s gender and race/ethnicity as well as each student’s university GPA upon entering the class. These data were obtained from the university administration and included before anonymizing the entire dataset several years ago. Since the numerical exam scores are given in a wide variety of ways by the many instructors of these classes, we have z-scored each exam in each course to give each exam grade distribution the same average, 0, and the same standard deviation, 1.
All together the database included 26,766 students with 16,485 identified as female students and 10,246 identified as male students. The exam data include 5,567 students appearing only once, 7,324 students appearing in only two of the classes and only once in each, and 13,089 students appearing exactly once in each of the three classes in the 3-course series. The three retake classes had a total of 397 students.
The number of individual exams given during the term varied from a minimum of 2 midterms given in each of seven classes to a maximum of 9 weekly exams given in each of 6 classes with the most common number given being 4 in-term exams in each of 111 classes (three of those are the retake classes). In addition, the databases from 203 classes also had final exam grades and so those classes had all exams accounted for.
III.1 Classifying the Stakes of an Assessment
For the purposes of this paper we have decided to classify the ‘stakes’ of an exam into three different categories. We use the same definition of high-stakes assessments as those typically defined in the literature. High-stakes assessments are timed exams that are usually in-person and allow limited use of resources. For example, a student might be able to bring a note card, but they would not be able to use the internet during the exam. Usually these assessments are compared to low-stakes assessments like labs and/or homework. In this paper we use a different definition of low-stakes. We consider the first try on an in-term exam that has a retake option to be a ‘low-stakes’ assessment because a second try is free to any student and could completely erase that first score. We also introduce the idea of a ‘medium stakes’ assessment. We consider the exams given as the term progresses to be ‘medium stakes’ both because each individual in-term exam counts for a much smaller portion of the course grade than the final exam and because usually the class syllabus described in-course grading rules telling students that the lowest in-term exam would be dropped before calculating the course grade. These definitions are outlined in Table 1.
Table 1: Descriptions of the magnitude of the stakes associated with an exam. The stakes associated with one exam can be thought of as determined by the minimum amount that the exams can possibly contribute to the course grade divided by the total number of exams
Min. possible
Exam Exam contribution
Stakes Type to course
grade by full
set of exams
Retake class - Set of
Low four first-attempts 0%
at short in-term exams
Non-retake class - Set
Medium of four to nine 40%
short in-term exams
or two midterm exams
Both types of class -
High One long final exam 40%
IV Results
IV.1 In-term exams in retake classes
In a previous paper [1] we showed that the gender gap in course grades was zero for a set of four classes that offered retake exams to their students. In a retake class students took an in-term exam, got their scores on that exam, and then (the next week) had a chance to take a different exam (the retake exam) that covered the same material. Students were told that if they scored higher on the retake then the retake score would supplant the grade on their first attempt but if they scored lower on the retake then the retake and the first attempt scores would be averaged to give the recorded exam score. In these classes, there were a total of four in-term exams and so four in-term retake exams. Here we take a closer look at the details of these first attempts and the exam scores after the subsequent retakes.
Our expectation was that students who had high scores on their first attempt would not risk lowering that score and so would not take a retake exam. We found that an average first attempt was followed by 34% of the students taking the retake a week later. The retake classes were anomalous in having zero gender gap in course grades (i.e. female students had the same average course grades as male students unlike in the regular courses) suggesting the following possible gap-closing scenario regarding retakes: i) female students may be more likely to take retake exams to improve their grade and/or ii) female students may make larger gains on retake exams. Examination of the data shows that the first possibility is wrong. We find that 78% $\pm$ 4% of male students retook at least one exam and 74% $\pm$ 3% of female students retook at least one exam. All together male students retook an average of 1.41 $\pm$ 0.09 exams and female students retook an average of 1.33 $\pm$ 0.07 exams. It seems that the two genders had about the same need for retake exams.
We can examine the second possibility mentioned above by measuring the gender gap on each exam and also the gender gap after each retake had been scored. As noted in section III, the grades from the 12 exams from the three different courses were normalized so that i) each set of first-attempt scores from each exam in each course has the same average, = 0, and standard deviation, = 1, and ii) the recorded grades (i.e. after retake) from each exam from each class were similarly z-scored. The result of this is that each gender gap will be measured in units of standard deviations of the exam scores for that exam. We will also control for each student’s incoming GPA in our estimates of each in-term exam gender gap. We do this because the exam-score gender gap in a class has a small but clear dependence on the GPA gender gap in that class (see Appendix A which shows these class-dependent gender gap data) and we wish to remove that dependence from our in-term exam results. We control for GPA using hierarchical linear modeling (HLM) with the students as lowest level in the hierarchy, the set of classes as the next higher level, and then the part of the 3-course series (A, B, or C) makes up highest level of the hierarchy. We use all three levels because our retake classes cover all three levels. We use the variables $GPA$ for a student’s GPA and $Female$ = 1 for female students and = 0 for male students and fit the following model eight times, once for each of the four first-attempt exam scores taken through the term and once for each of the four exam scores recorded after the retake scores are factored in,
$InTrmExm=b_{0}+b_{GPA}GPA+b_{Female}Female$ (1)
so $b_{Female}$ is the (GPA-controlled) gender gap for each fit of in-term exam scores. We also average each student’s first-attempt exam scores over the term and fit our model to those and we average each student’s after-retake scores and fit our model to those. So $InTrmExm$ can be one of the four first-attempt exam results, one of the four after-retake exam results, or one of the averages over the full term.
Figure 1 shows these gender gaps for each first attempt and also after each retake exam. These results are plotted as a function of the week the exam was taken through the 10-week term. The overall averages, after all four in-term first-attempts and their retakes were graded, are shown to the right of the end of the term. Notice that students identifying as female outscored students identifying as male on three of the four first-attempt in-term exams. And the average of all first-attempt exams is slightly positive (female students out-performing male students) but the average of all in-term exams after retakes are scored is slightly negative (retake exam scores benefited male students slightly more than female students). So the idea that the retakes allowed female students to improve their first attempt scores in a way that reduced the gender gap is also untrue. This result suggests a third possible effect of allowing retake exams. iii) The knowledge that retakes are allowed lowers the stakes of each first-attempt from what it would be without retakes. So it may be that the scores are gender neutral because the stakes of the first-attempt exams are low. Specifically, we suggest that the first-attempt has lower stakes in a retake class than it does in the regular courses where the first attempt is also the last attempt. In Figure 1 we have also included the week-by-week GPA-controlled in-term exam gender gaps for three sets of non-retake classes for which the week-by-week exam pattern through the term is known. In the rest of the paper we use the ideas of low-stakes, medium-stakes, and high-stakes exams in comparing the retake classes with the 254 non-retake classes for which we have exam data.
Figure 1: As explained in the text, each in-term exam was z-scored on a class-by-class basis and the gender gap for each group of classes was determined with HLM while controlling for each student’s incoming GPA. This in-term exam score gender gap is plotted as a function of time through the 10-week term at the times the exams and retakes were taken. On the right, after the end of the term, are the overall averages for those first-attempt exams and for the recorded scores after retakes were completed. One sees that the female students generally had higher exam scores (i.e. a positive gender gap) after the first attempt and that, overall, retakes favored the male students. For reference the gender gaps for 179 non-retake classes are also shown on a week-by-week basis through their terms.
IV.2 Low-stakes versus medium-stakes assessments
Instructors in this series of classes determined course grades almost completely by averaging the exam grades using a weighting of their choice. A weighting that was common (almost universal) was to 1) drop the lowest score among the in-term exams (either one in-term exam or one-half of one midterm) and then 2) weigh the final exam somewhere between 40% and 60% depending on whether the in-term exam average was higher or lower, respectively, than the final exam score. The most common numbers of in-term exams were four and eight so the grade weight of a single in-term exam ranged from roughly 5% to 15% of the course grade. We consider these exams to be medium-stakes assessments. The final exam carries several times the weight of any particular in-term exam so we consider it to be a high stakes assessment. For the retake courses we consider the first-attempt of every exam set that students take to have low stakes because the student knows that they are free to take another exam and discount the first one if they choose to.
We are interested in comparing the gender gap resulting from low-stakes in-term exams with that from medium-stakes in-term exams. For this reason we compare the gender gap in the first-attempt exams from the retake classes with the gender gaps in the 254 non-retake classes in our dataset. Since different classes had different numbers of in-term exams we will just average the z-scored in-term-exam scores for the entire term for each student in each class. Also remember that the gender gap for in-term exams in a class has a small but clear dependence on the gender gap for GPAs in that class so we will control for GPA in our analyses.
We model the comparison of the 3 retake classes with the 254 regular classes using the same kind of 3-level hierarchical regression analysis that was previously discussed. Again, the student level is the lowest level, then students are grouped in classes (2nd level), and the classes are grouped within the course series (A, B, and C in the series identify the highest level). We use all three levels because our database covers all three levels and also because the different classes may have slightly different grade penalties for female students. We again use the variables $GPA$ for a student’s GPA, $Female$ = 1 for female students and = 0 for male students, and now include $Retake$ = 1 for the three retake classes and = 0 for the 254 regular classes. The model we fit for the in-term exam z-scores is
$\begin{split}InTrmExm&=b_{0}+b_{GPA}GPA\\ &+b_{Female}Female+b_{Retake}Retake\\ &+b_{Female*Retake}Female*Retake\end{split}$ (2)
where the coefficient $b_{Female}$ represents the GPA-controlled gender gap and the coefficient $b_{Female*Retake}$ measures the difference in that gender gap when retakes are possible.
The coefficients resulting from this HLM fit are given in Table 2. Recall that we consider the first-attempt exams we are using for the retake classes to be low stakes. In the 254 non-retake classes which had medium-stakes in-term exams the average grade penalty ($b_{Female}$) that female students receive over the term is about 0.23 standard deviations (of the exam average distribution). The interaction coefficient ($b_{Female*Retake}$) for classes with low-stakes exams also has magnitude 0.23 but is positive so that the low-stakes exams net gender gap is, as expected, approximately zero. Seeing a gender gap disappear when the only change seems to be the stakes of the exams might cause one to pause before claiming that the gender gaps in the medium-stakes exams are actually measuring a gap in physics knowledge/ability.
Note that even though the coefficient for $Retake$ is negative, indicating (perhaps) that men do worse under retake conditions, this is only because the grades are z-scored. In our previous paper [1] we showed that, on average, both genders had higher grades in the retake classes. Also, as shown in Appendix C, the gender gap decreases in size in retake classses for each racial/ethnic group which are sufficiently represented in our data. Finally, as discussed in Appendix B, to decide if these coefficients or our resulting conclusions are affected by the fact that many students are in our database more than once (because they have taken more than one of the classes in this three-term series) we have tested the effects of limiting each student to just one appearance in the database. We find that the coefficients and our resulting conclusions are unchanged.
Table 2: The coefficients from an HLM fit to Eq. 2 are shown along with their standard errors, z-statistics, and p-values. Included are 257 classes that had a total of 57,806 students whose GPA’s are known. The interaction term suggests that the gender gap is significantly different (reduced) for the retake classes.
Coeff. Value Error z-statistic p-value
$b_{GPA}$ 0.8121 0.0047 173.0 $<10^{-3}$
$b_{Female}$0.2262 0.0046 -49.0 $<10^{-3}$
$b_{Retake}$0.169 0.052 -3.24 0.001
$b_{Female*Retake}$ 0.234 0.057 4.10 $<10^{-3}$
$b_{0}$2.310 0.018 -128.1 $<10^{-3}$
IV.3 Highest stakes assessments
Since 203 of the classes in our database have final exam scores it seems useful to examine those data because the final exam scores contribute much more to the course grade than any in-term exam and so can be considered to be high-stakes exams. Unfortunately we don’t have a low-stakes version of the final exam to compare with the high-stakes versions. However, we do have the medium-stakes in-term exam scores in each class which should be good predictors of final exam scores in that class because they are written and graded by the same instructors and largely cover the same material. In fact, we find a moderate/strong correlation between the in-term exam average and the final exam score (correlation coefficient = 0.65, $p<10^{-3}$) but a smaller correlation between student GPA and either in-term or final exam score (both correlation coefficients are about 0.56, $p<10^{-3}$). We’ve already suggested that the medium-stakes in-term exam scores include a bias against female students, so it may be that using these scores as predictors of final exam scores will not only control for the demonstrable knowledge/ability of each student in the particular course but may also control for any gender bias inherent in exams. If this were the case then the controlled final exam gender gap should be approximately zero. To test this idea we will use HLM to fit the following model
$\begin{split}FinalExam&=b_{0}+b_{InTrmExm}InTrmExm\\ &+b_{Female}Female+b_{Retake}Retake\\ &+b_{Female*Retake}Female*Retake\end{split}$ (3)
where we include the variable $Retake$ and the interaction term even though the retake courses had the same sort of high-stakes final exam as all the other courses. If the gender bias on in-term exams is roughly the same as the gender bias on final exams then the coefficient $b_{Female}$ from this fit should be nearly zero.
The resulting coefficients from the HLM fit to Eq. 3 are given in Table 3. The most notable coefficient is $b_{Female}$ and the most notable thing about this coefficient is that it is quite clearly not zero. In fact, the (in-term exam controlled) final exam gender gap is half the size of the gender gap for the in-term exams alone. This tells us that the full gender gap on the high-stakes final exam is about 50% larger than that of the medium-stakes in-term exams. So the high-stakes final exam has a much larger gender gap than the medium-stakes in-term exams which, themselves, have a much larger gender gap than the low-stakes first-attempt in-term exams in the retake classes. And $b_{Female*Retake}$ is consistent with there being a similar (in-term-exam controlled) gender gap in the high-stakes final exam in the retake classes. Again, it seems likely that the importance (or weight) of the exam (the exam stakes) is closely related to the size of the grade penalty received by female students.
Table 3: The coefficients from an HLM fit to final exam z-scores (equation 3) are shown along with their standard errors, z-statistics, and p-values. Included are 203 classes that had a total of 49,085 students. The interaction term suggests that the gender gap is not significantly different for the retake classes.
Coeff. Value Error z-statistic p-value
$b_{InTrmExm}$ 0.9656 0.0052 187.4 $<10^{-3}$
$b_{Female}$0.1155 0.0071 -16.2 $<10^{-3}$
$b_{Retake}$ 0.025 0.066 0.38 0.705
$b_{Female*Retake}$0.034 0.080 -0.42 0.676
$b_{0}$ 0.0744 0.0056 13.23 $<10^{-3}$
V Discussion
In our introduction we present many student-level variables that may contribute to the gender gap in introductory physics. But these factors; preparation, interest, anxiety, stereotypy threat, self-efficacy, identity, and belonging are all deeply intertwined with both larger influences of society, and the cultural context of the physics discipline. For instance, results from Leslie et. al. [24] on the gender imbalance in academic fields (including physics) that are felt to require “innate talent” rather than persistent effort has led these researchers as well as others (Muradoglu et. al. [25]) to argue that, in fields perceived to require “innate talent” women (and other groups that are often intellectually marginalized) have more negative experiences, feel less welcome, have more “imposter” feelings, and are more vulnerable to stereotype threat [26]. This general idea can connect the various gender gap explanations that we have listed above. We postulate that many of these feelings of otherness are exaggerated in a high-pressure environment. With this stated assumption we can then argue that the reason the gender gap exists is not because of all these negative biases against women, it’s because these biases exist AND we put women in a high pressure environment.
Furthermore, an extremely important takeaway from this logic is that tests are not measuring what we think they are measuring because if they were, performance gaps between demographic groups would not depend on the stakes of the particular exams as Sec. IV.3 shows they do. The implications are that if we care about measuring physics knowledge and skills, we need to be looking at ways to reduce or eliminate the stakes associated with our assessments because since the stakes of a given timed in-person assessment correlate with the size of the gender gap, exams are currently not good measures of physics knowledge and skills across demographic groups.
Another problem with a focus on the individual student level issues that large groups of students may be having in physics is that this leads too naturally to a consideration of those groups of students as deficient or inadequate in some way. Researchers tend to use a Student Deficit model rather than the idea that these students are just reacting, on average, the way anyone would if they were part of a group stereotyped by society. Indeed, results from Salehi et al [8] fit this picture - they find that the gender gaps for exams are much larger in general science and engineering courses where women are less represented compared to biological science courses where women are more represented and may therefore feel more like they belong in that space.
Cotner and Ballen [2] cut the Gordian knot made up of interconnected possible student deficits by suggesting the easiest thing to do is to simply change the structure of the course. They proposed using a Course Deficit model rather than a Student Deficit model. Our recent work [1, 27] supports the idea that changing the course structure is enough to eliminate equity gaps. While studies that examine the effects that society (and more specifically disciplinary cultures) have on different demographic groups are important and necessary, we should not present those effects as the reasons for the differences between the performances of demographic groups. If we move in that direction we miss the forest for the trees. For example, the solution to the gender gap in physics exams isn’t to make women feel less anxious in a stressful environment, it’s to change the environment to be less stressful. This is the power that we have as instructors. We might not be able to change the impact society has on women in physics, but we can mitigate the effects of those impacts in our courses.
Instructors reading this paper might consider adoption of retake exams or other assessments that lower the stakes associated with testing in introductory physics courses.
VI Conclusions
Building from our prior research on this dataset, [1] where we found that retake exams were associated with a zero gender gap in introductory courses, we rule out the possibility that this happens because women are more likely to take advantage of the retake opportunity, as men are just as likely to take advantage of the retakes. We also find that the reason for eliminated gender gap doesn’t appear to be that women need an extra chance to fail and catch up (presumably due to any differences in incoming physics experiences) because women appear to do as well on the first try of a retake exam as the men.
Instead we find that women do as well as and perhaps slightly better (though not significantly so) than men on the first try of a retake than their male peers. We also find that women do significantly better overall on all of the exams (independently true for both first and second tries) in the retake course than they do in the courses without retake exams. Finally, if the stakes of the final exam didn’t matter for the gender gap, we would find that controlling for in-term exam scores (an assessment of the same type, but not the same stakes) would zero the gender gap on the final exams. Instead we find that controlling for in-term exam scores on the final exams in courses without retakes, does not remove the gender gap in final exams, again suggesting that the stakes of the exam are responsible for the gender gap in physics.
Appendix A Detailed retake exam data
In the text we have noted that the class gender gap for in-term exams has a clear dependence on the GPA gender gap in that class. This is shown in Fig. 2 where the class-average in-term exam gender gap is plotted as a function of that class’ GPA gender gap for all 254 regular classes. We have also grouped the first-attempts of the three retake classes together and plotted them as a single point so that one can see how an average retake class might compare to the non-retake classes. In previous work we have noted that there seems to be a large exam/grading noise that is not easily explained by differences in student GPAs. Thus, class-dependent gender gaps, which involve differences in noisy grade distributions might be expected to be even noisier. The standard deviation, over classes, of the gender gap is about half of the average gender gap, showing the large variety of results even in classes that have the same instructional materials. Nevertheless, the average retake class sits at the upper edge of the distribution over classes.
Figure 2: In-term-exam average z-score for female students minus the same for male students plotted as a function of average GPA of female students minus the same for male students for 257 classes in the intro-physics for bio-scientists series at one R1 university. The three retake classes for which we have data on the first-attempt exam scores were computed together to show the average effect upon taking an exam for the first time when a retake is possible. No retakes were possible in the other classes so that those exams had higher stakes.
Appendix B Including a student only once doesn’t change our numbers much
As we point out in Sec. III, many of the students in our dataset are represented more than once because they were enrolled in, and received grades in, more than one class included in the data. One worry we had was that including these students more than once might influence our numbers or our conclusions. To check this we have, for each of the 26,766 unique students in the datafile, randomly chosen only one of their classes to include in an analysis. We made 10 of these random choices in order to check the coefficients and errors in Table 2. We find that the model coefficients we present in the text do not change significantly and that the error estimates change as $\sqrt{1/N}$ , where $N$ is the total number of sets of student grades, as one expects if one simply had fewer data points. So the results and conclusions we report in the text are essentially unchanged by having a set of students present in the data more than once because they took more than one class in the series.
Table 4: The average coefficients from 10 HLM fits to equation 2 (where each fit corresponds to a different random choice of which one class, for each student, is included) are shown along with their standard errors, z-statistics, and p-values. The comparison with Table 2, which includes all classes of all students, shows that the only important effect is to increase the error estimates by the square root of ratio of number of students included in the fit. For this table the average number of students included (because we know their GPA’s) averages 24,641. And the retake classes include only 368 students instead of the 57,806 students and 396 students, respectively, included in Table 2.
Coeff. Value Error z-statistic p-value
$b_{GPA}$ 0.8052 0.0072 112.0 $<10^{-3}$
$b_{Female}$0.2226 0.0072 -31.0 $<10^{-3}$
$b_{Retake}$0.164 0.055 -3.00 0.003
$b_{Female*Retake}$ 0.233 0.061 3.85 $<10^{-3}$
Appendix C Results from the intersections of gender and race/ethnicity
In our original paper [1] we pointed out that the classes offering retake exams did not have statistically significantly smaller grade gaps between racial/ethnic groups historically marginalized within physics and their peers from groups that have not been marginalized. In the interests of completeness in this paper we address the intersection of gender with race/ethnicity. We again use Eq. 2 to find the GPA-controlled gender gap difference between the retake and non-retake classes but now we show the results of 12 separate fits to that model, one for each of the racial/ethnic groups in our database for which both genders are populated in both non-retake and retake classes.
In Fig. 3 we show the results for $b_{Female*Retake}$, as determined from a fit to Eq. 2 for each fit to the particular racial/ethnic group described in Table 5. One sees that these differences are all positive although most of them are not significantly different than zero. However, it is notable that several are well above zero.
Figure 3: The, GPA-controlled, gender gap differences between the first-attempt in-term exam average in the retake classes and in-term exams average in the non-retake classes. These are shown for each separate racial/ethnic group for which there are both male students and female students identifying with that racial/ethnic group in both the retake classes and the non-retake classes. We are actually plotting $b_{Female*Retake}$, as determined from a fit to Eq. 2 for each fit to the particular racial/ethnic group described in Table 5.
Table 5: The ethnicities in Fig. 3, as described in the data obtained from the administration of the R1 university of this paper.
Symbol Ethnicity
AF African-American/Black
CH Chinese-American/Chinese
EI East Indian/Pakistani
FP Filipino/Filipino-American
IA Indigenous American/American Indian/
Native American
JA Japanese-American/Japanese
KO Korean-American/Korean
LA Latino/Other Spanish
MX Mexican-American/Mexican/
Chicano
OA Other Asian-American/Other Asian
VT Vietnamese-American/Vietnamese
WH White/Caucasian
Appendix D Grade scale effect on gender gap
We haven’t discussed issues around grade scale and gender in earlier papers so we will do some of that here. The retake course exams discussed in this paper were graded on a very coarse grade scale. Specifically, an answer that had the physics explained (including setting up equations) fully and correctly received one point (minor arithmetic errors were ignored), an answer that had the physics almost, but not perfectly, complete and correct received 2/3 of a point, and every other answer received zero points. The one-point answers would have received A+ or A or A- under the regular grading in the non-retake classes and the 2/3-point answers would have received B+ or B or (maybe) B- in the non-retake classes. One might wonder if this kind of coarse grade scale used in the retake classes could explain the lack of a gender gap in the retake classes.
Even though the regular non-retake classes used a variety of grade scales, most of these classes used either a 4-point scale or a percent scale so, for these classes, we can identify the fraction of A-grades or B-grades on specific in-term-exam items received by female students and compare to the fraction received by male students. For this subset of 156 non-retake classes, we find that female students received A-grades on 40.02% $\pm$ 1.3% of their individual exam answers and B-grades on 15.24% $\pm$ 0.07%. On these same exams male students received A-grades on 45.05% $\pm$ 1.7% of their exam answers and B-grades on 15.03% $\pm$ 0.07%. So female students averaged about 5% fewer A’s and about the same number of B’s as male students. If we take these grades from the non-retake classes and rescore all A-grades to be one point, all the B-grades to be 0.67 points, and the rest of the grades to be zero, then we find the standard deviation of the resulting re-scored in-term exam grade distribution is about 0.2. So the average normalized gender gap under this kind of coarse grading would be about (average difference of 0.05 fewer A’s) multiplied by (1 point per A) and divided by (standard deviation of 0.2) = 0.25 which, given the crudeness of these grade changes, is reasonably the same as we saw for $b_{Female}$ in Table 2. We suggest that the coarse grading had a negligible effect on the gender gap.